Circular No. 6730

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 A. J. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofísica Es-

pacial y Física Fundamental, Madrid; J. Iglesias, L. M. Cairos, J. Vilchez, A. Mora, C. Gutiérrez, J. Licandro, and V. Bejar, Instituto de Astrofísica de Canarias (IAC), La Laguna; and J. Greiner, Astrophysikalisches Institut, Potsdam, report: “Following the detection of the x-ray afterglow from GRB 970828 by RXTE (IAUC 6726, 6727), R-band images were obtained (on Aug. 29.03 and 29.95 UT) with the Teide Observatory 0.8-m IAC-80 telescope, and unfiltered (Aug. 29.91, 30.09, and 31.10) and B-band (Aug. 30.92) images were obtained with the Calar Alto 2.2-m telescope (+ CAFOS). About 90 objects are detected in the ASCA error box (IAUC 6729), down to mag 23. No variations larger than 0.3 mag are seen for objects brighter than mag 22.5. An object of magnitude B = 22.7 ± 0.3 is clearly seen in the 25-min B-band image but only barely visible in the 60min, co-added unfiltered image. Its position (± 1′′) is α = 18^{h}08^{m}32^{s}.7, δ = +59^{o}18′22′′.9 (equinox 2000.0) — i.e., 16′′ west and 13′′ north of a galaxy (B = 20.2) seen on the Palomar Sky Survey. It lies in the intersection between the RXTE and ASCA error boxes.”

D. A. Frail, National Radio Astronomy Observatory and S. R. Kulkarni, California Institute of Technology, communicate: “The VLA was used to observe the field of GRB 970828 (IAUC 6726) on three occasions. A single radio source was detected within the area defined by the intersection of the refined RXTE ASM error ellipse (IAUC 6728) of GRB 970828 and the ASCA error circle (IAUC 6729). The position of the radio source is α = 18^{h}08^{m}25^{s}.51 ± 0^{s}.05, δ = +59^{o}18′08′′.5 ± 0′′.4 (equinox 2000.0). On Aug. 31.19 and Sept. 1.27 UT, the source had a mean flux of 0.36 and 0.27 mJy, respectively, at 6 and 3.6 cm; the rms error in the flux is 0.03 mJy. There is some evidence of variability, since this radio source was not detected at 6 cm on Aug. 28.88. A 4σ upper limit to the flux of the source at this epoch is 0.18 mJy.”

S. G. Djorgovski and S. C. Odewahn, California Institute of Technology; and Kulkarni and Frail further report: “CCD images of the radio field noted by Frail and Kulkarni (above) were obtained by C. C. Steidel, K. L. Adelberger, and M. Kellog at the prime focus of the 5-m Hale telescope on Aug. 30, 31, and Sept. 1 UT. We have identified a faint object (magnitude R ∼ 24.5 ± 0.5) that is coincident with this radio source. The source did not vary significantly over the three-night interval.”

1997 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green